# PhyGDPO: Physics-Aware Groupwise Direct Preference Optimization for Physically Consistent Text-to-Video Generation

### v2

@@ Abstract
Recent advances in text-to-video (T2V) generation have achieved good visual quality, yet synthesizing videos that faithfully follow physical laws remains an open challenge. Existing methods mainly based on graphics or prompt extension struggle to generalize beyond simple simulated environments or learn implicit physical reasoning. The scarcity of training data with rich physics interactions and phenomena is also a problem. In this paper, we first introduce a Physics-Augmented video data construction Pipeline, PhyAugPipe, that leverages a vision–language model (VLM) with chain-of-thought reasoning to collect a large-scale training dataset, PhyVidGen-135K. Then we formulate a principled Physics-aware Groupwise Direct Preference Optimization, PhyGDPO, framework that builds upon the groupwise Plackett–Luce probabilistic model to capture holistic preferences beyond pairwise comparisons. In PhyGDPO, we design a Physics-Guided Rewarding (PGR) scheme that embeds VLM-based physics rewards to steer optimization toward physical consistency. We also propose a LoRA-Switch Reference (LoRA-SR) scheme that eliminates memory-heavy reference duplication for efficient training. Comprehensive experiments show that our method significantly outperforms state-of-the-art open-source methods on the PhyGenBench and VideoPhy2 datasets.
@@ 1 Introduction
The emergence of direct preference optimization [49] (DPO) may provide a potential solution to address the limitations of SFT. However, directly using DPO may encounter three problems: (i) Lack of training data pairs that comprehensively capture physical activities, interactions, and phenomena. (ii) The supervision is mainly based on the Bradley–Terry (BT) probabilistic model, which only compares pairwise data samples. Such binary win/lose feedback lacks the guidance of physics-aware rewarding and struggles to capture inherently holistic global preference signals, such as physical plausibility and motion smoothness. (iii) Vanilla DPO duplicates the training model as a fixed reference, which consumes substantial GPU memory and significantly decreases computational efficiency.
@@
To address these issues, we firstly propose a data construction pipeline, PhyAugPipe, that exploits a vision-language model (VLM) to filter text-video data pairs capturing rich physical interactions and phenomena from a large T2V data pool by parsing the entities and reasoning their actions and outcomes using our designed chain-of-thought (CoT) [60] rule for each data sample. Then we use PhyAugPipe to collect a training dataset, PhyVidGen-135K, containing over 135K text-video data pairs. Secondly, we formulate a principled Physics-aware Groupwise Direct Preference Optimization (PhyGDPO) framework for physically consistent T2V generation. Our PhyGDPO is based on the groupwise Plackett-Luce (PL) probabilistic model, which stimulates the probability distribution over a ranked list of candidate videos and thus enables holistic preference adaptation rather than being constrained to isolated pairwise judgments. Plus, the original DPO training uses the model itself as reward [49] but the model itself can not provide sufficiently good signals and knowledge of physics. Hence, we propose to embed more physics reward during training. Unlike GRPO [51], which suffers from slow speed and unstable optimization, we propose a Physics-Guided Rewarding (PGR) scheme that leverages a physics-aware VLM to guide the DPO data sampling favor more challenging physical actions and allow physics-violating samples to exert stronger influence during preference optimization. To improve the DPO training efficiency and stability, we also propose a LoRA [26]-Switch Reference (LoRA) mechanism that does not need to copy the full model as the reference, which occupies redundant GPU memory. Instead, we attach a trainable LoRA module with a switch to the frozen backbone. Benefit from our dataset and proposed techniques, PhyGDPO significantly improves the physical plausibility of the T2V model, Wan2.1-T2V-14B [55], and yields better results than the two recent best commercial T2V models, OpenAI Sora2 [6] and Google DeepMind Veo3.1 [18] on some challenging action categories, as shown in Fig. 1, 4, 5, and 6.
Our contributions can be summarized as follows:
@@
  • We formulate a principled DPO framework, PhyGDPO, for physically consistent text-to-video generation.
@@
  • We design PGR to guide data sampling and preference optimization favor challenging physical actions. We propose LoRA-SR scheme to reduce GPU memory occupancy for more efficient and stable DPO training.
@@ 2.1 Physics-Augmented Video Data Construction
Our physics augmented video data construction pipeline (PhyAugPipe) is shown in Fig. 2. We use a VLM, Qwen-2.5-72B-Instruct [65], with our chain-of-thoughts (CoT) in (b) to filter videos with rich physics interactions and phenomena in (c) from a T2V data pool in (a). Then in (d), we cluster the filtered data according to their action categories and sample the data with the rewards of physics-aware VLM, VideoCon-Physics [3], in (e).
@@
Figure 3: Overview of our PhyGDPO framework. The text prompts without physics reasoning extension in (a) are fed into the T2V model. We propose a LoRA-switch reference scheme in (b) to save the GPU memory and increase the training stability. In (c), PhyGDPO is based on the groupwise Plackett-Luce (PL) probabilistic model and adopts a physics-aware VLM, VideoCon-Physics [3], to reward the DPO training with the real video as the wining sample as it has perfect physics.
Data Sampling with Physics Rewarding. As different action categories vary significantly in physical complexity thus leading to different generation difficulties for the T2V model, we propose to balance the training data accordingly. In particular, we first evaluate how well the T2V model performs on each action category. Based on the performance distribution, we adjust the sampling ratio by allocating more samples to the categories where the model performs poorly. As shown in Fig. 2 (d), we first select the top-$n_{c}$ samples with the highest semantics matching scores within each category as its representatives. Then we employ a physics-aware VLM, VideoCon-Physics [3], to evaluate each video representative. VideoCon-Physics outputs a semantics adherence score and a physics commonsense score $\in[0,1]$ to measure the physics plaucibility of the video. We average the two scores to obtain the overall score for each video. Then we compute the mean score of all representatives within each category to obtain the category score. In this way, we derive the score distribution histogram $\mathcal{S}_{f}$. Then we sample the data according to the difficulty of each action category. More specifically, we define the difficulty of the $k$-th action category as $d_{k}=1-\mathcal{S}_{f}(k)$ and assign a sampling weight following an exponential form as $r_{k}=\text{exp}(\tau d_{k})$, where $\tau$ is a hyperparameter that controls how strongly the sampling favors difficult action categories. Given the total sampling budget $N$, the number of data pairs sampled from the $k$-th action category is formulated as
$\normalsize\mathcal{H}_{r}(k)=\min\bigg(\mathcal{H}_{f}(k),\,N\cdot\frac{r_{k}}{\sum_{j=1}^{K_{a}}r_{j}}\bigg),$ (1)
where $\mathcal{H}_{r}$ denotes the distribution histogram after data sampling with physics rewarding. This sampling strategy allocates more data to challenging actions, encouraging the model to learn more complex physics during training.
@@ 2.2 Physics-Aware Groupwise Direct Preference Optimization
Existing DPO algorithms are based on Bradley–Terry (BT) probabilistic model, which compares pairwise data samples, showing limitations in capturing global and holistic preference signals and aligning with human feedback. Besides, previous DPO [54, 36] mainly use the T2V model itself as the reward [49], but the model itself exhibits weak physics reasoning ability as it usually generates physically inconsistent videos. To address these issues, we formulate a Physics-Aware Groupwise Direct Preference Optimization (PhyGDPO), as shown in Fig. 3.
Groupwise Probability. We denote the reward as $r(c,x_{0})$ with the generation $x_{0}$ and condition $c$. Different from the normal DPO, we start from the groupwise Plackett-Luce (PL) probabilistic model, as shown in Fig. 3 (c). We adopt the real video as the winning case $x_{0}^{w}$ because it always follows physical laws and a set of generated videos as the losing cases $\mathcal{G}^{l}(c)=\{x_{0}^{l_{1}},\ldots,x_{0}^{l_{m}}\}$. The preference probability of the PL model is formulated as
@@
$\normalsize\mathcal{L}_{\mathrm{PL}}(\phi)=-\,\mathbb{E}_{c,\,\mathcal{G}^{l}(c)}\!\Bigg[\log\frac{\exp\!\big(r_{\phi}(c,x^{w}_{0})\big)}{\exp\!\big(r(c,x^{w}_{0})\big)+\sum_{j=1}^{m}\exp\!\big(r_{\phi}(c,x^{l_{j}}_{0})\big)}\Bigg].$ (3)
@@
$\normalsize\log\frac{p_{\theta}(x_{0}|c)}{p_{\psi}(x_{0}|c)}=\sum_{k=1}^{T}\log\frac{p_{\theta}(x_{t_{k-1}}\mid x_{t_{k}},c)}{p_{\psi}(x_{t_{k-1}}\mid x_{t_{k}},c)}\;\triangleq\;\sum_{k=1}^{T}\Delta_{k}.$ (6)
@@
Please refer to the supplementary for the detailed proof process of this inequality. Subsequently, we can further formulate the upper bound of $\mathcal{L}_{\text{GDPO}}(\theta)$ into a single sample within each group for efficient training as
@@
$=\mathbb{E}_{c,\mathcal{G}^{l}(c),k,j}\Big[-\gamma_{j}\log\sigma\big(-\alpha_{j}\beta T(\Delta_{k}^{w}-\Delta_{k}^{l_{j}})\big)\Big].$
@@
Physics-Guided Rewarding. As aforementioned, normal DPO framework uses the T2V model itself as the reward like Eq. (4), which cannot provide sufficiently good physics knowledge. To embed more physics guidance into our DPO framework, we parameterize $\gamma_{j}$ and $\alpha_{j}$ in Eq. (9) by the normalized semantics adherence score $s^{sa}_{j}$ and physics commonsense score $s^{pc}_{j}$ predicted by a physics-aware VLM, VideoCon-Physics [3], as
@@
$\alpha_{j}$ $=\alpha_{\text{min}}+(1-\alpha_{\text{min}})\cdot\tanh\big(\kappa_{\alpha}(v_{j}-b_{\alpha})\big),$
where $v_{j}$ measures the physics difficulty and mainly modulates $\gamma_{j}$ and $\alpha_{j}$, $\alpha_{\text{min}}$ sets a lower bound to avoid vanishing gradients, $\tanh(\cdot)$ provides a smooth nonlinear mapping that gradually sharpens the preference decision boundary as the gap increases, and the sigmoid function $\sigma(\cdot)$ ensures a bounded and smooth adjustment of $\gamma_{j}$, stabilizing the optimization. $\alpha_{j}$ adaptively adjusts the sharpness of the preference comparison, while $\gamma_{j}$ amplifies the physics-guided reward for samples with lower physical plausibility. They are controlled by the hyperparameters $\lambda$, $\kappa_{\alpha,\gamma}$, $b_{\alpha,\gamma}$, and $\alpha_{\text{min}}$. Our designed physics rewards $(\alpha_{j},\gamma_{j})$ balances the training stability with adaptivity, allowing physics-violating samples to exert stronger influence during optimization.
@@
$\log\frac{p_{\theta}(x_{t-h}\mid x_{t},c)}{p_{\psi}(x_{t-h}\mid x_{t},c)}$ $\approx-\frac{h}{2\varepsilon}\Big(\ell_{\theta}(x_{t},t)-\ell_{\psi}(x_{t},t)\Big),$ (13)
@@
$\normalsize\mathcal{L}=\mathbb{E}_{c,\mathcal{G}^{l}(c),k,j}\Big[-\gamma_{j}\log\sigma\!\big(-\alpha_{j}\beta T[(\ell_{\theta}^{w}-\ell_{\psi}^{w})-(\ell_{\theta}^{l_{j}}-\ell_{\psi}^{l_{j}})]\big)\Big],\vskip 2.84526pt$ (14)
where any global constant scale factor such as $h/2\varepsilon$ in Eq. (13) can be absorbed into the hyperparameter $\beta$. For simplicity, we denote $\ell_{\theta}^{w}=\ell_{\theta}(x_{t_{k}}^{w}\mid t_{k})$, $\ell_{\psi}^{w}=\ell_{\psi}(x_{t_{k}}^{w}\mid t_{k})$, $\ell_{\theta}^{l_{j}}=\ell_{\theta}(x_{t_{k}}^{l_{j}}\mid t_{k})$, and $\ell_{\psi}^{l_{j}}=\ell_{\psi}(x_{t_{k}}^{l_{j}}\mid t_{k})$ in Eq. (14).
LoRA-Switch Reference. Previous DPO methods usually copy the full model and fix it as the reference, which occupies redundant GPU memory and degrades computational efficiency. In addition, this full-copy strategy often leads to unstable and less effective DPO training because this strategy updates the entire set of the backbone weights and may cause the trainable model to quickly and dramatically deviate from the frozen reference model.
@@
To address this issue, we design a LoRA-Switch Reference (LoRA-SR) scheme. As shown in Fig. 3 (b), we freeze the backbone as the reference model $\psi$ and attach trainable LoRA modules to $\psi$ as the trained model $\theta$. Our LoRA-SR enables the reference and trainable models to share the same heavy backbone while flexibly switching the lightweight LoRA parameters for the training and reference modes. Thus, we can evaluate the trainable and reference terms without storing or loading an extra full model. This significantly increases computational efficiency and prevents the model from drifting too far from the reference, thus stabilizing the training process.

### v3

@@ Abstract
Recent advances in text-to-video (T2V) generation have achieved good visual quality, yet synthesizing videos that faithfully follow physical laws remains an open challenge. Existing methods mainly based on graphics or prompt extension struggle to generalize beyond simple simulated environments or learn implicit physics reasoning. The scarcity of training data with rich physics interactions and phenomena is also a problem. In this paper, we first introduce a Physics-Augmented video data construction Pipeline, PhyAugPipe, that leverages a vision–language model (VLM) with chain-of-thought reasoning to collect a training dataset, PhyVidGen-135K. Then we formulate a principled Physics-aware Groupwise Direct Preference Optimization, PhyGDPO, framework that uses real-world video as winning case to guarantee correct physics learning and builds upon the groupwise Plackett–Luce probabilistic model to capture holistic preferences beyond pairwise comparisons. In PhyGDPO, we design a Physics-Guided Rewarding (PGR) scheme that leverages VLM-based physical rewards to direct the optimization to focus on challenging physics cases. In addition, we propose a LoRA-Switch Reference (LoRA-SR) scheme that avoids full-model duplication as reference for efficient DPO training. Comprehensive experiments show that our method outperforms state-of-the-art methods on the PhyGenBench and VideoPhy2 datasets.
@@ 1 Introduction
The emergence of direct preference optimization [53] (DPO) may provide a potential solution. However, directly using DPO may encounter three problems: (i) Lack of training data pairs that comprehensively capture physical activities, interactions, and phenomena. (ii) The correct physics preference optimization cannot be guaranteed because vanilla DPO uses generated video as winning case and the physical realism of the generated video is limited and some challenging cases remain difficult to generate with correct physics. The supervision is mainly based on the Bradley–Terry (BT) probabilistic model, which only compares a single pair of generated samples. Such single binary comparison struggles to capture inherently holistic global preference signals of physical plausibility. (iii) Vanilla DPO copies the full model as reference, which consumes substantial GPU memory and decreases efficiency.
@@
To address these issues, we firstly propose a data construction pipeline, PhyAugPipe, that exploits a vision-language model (VLM) to filter text-video data pairs capturing rich physical interactions and phenomena from a large T2V data pool by parsing the entities and reasoning their actions with our designed chain-of-thought (CoT) [64] rule. We use PhyAugPipe to collect a training dataset, PhyVidGen-135K, containing 135K text-video pairs. Secondly, we formulate a novel Physics-aware Groupwise Direct Preference Optimization (PhyGDPO) framework for physically consistent T2V generation. PhyGDPO uses the real-world video that always follows physical laws as the winning case to guarantee correct physics learning. Different from vanilla pairwise DPO, PhyGDPO is based on the groupwise Plackett–Luce (PL) model, which captures the probability distribution over a group of candidate videos, enabling holistic preference adaptation beyond isolated pairwise comparisons. To further improve physics preference optimization, we propose a Physics-Guided Rewarding (PGR) scheme. PGR leverages a physics-aware VLM to guide data sampling and DPO training to focus on challenging actions and allows physics-violating samples to exert stronger influence. To improve DPO training efficiency and stability, we also propose a LoRA [28]-Switch Reference (LoRA-SR) scheme that does not need to copy the full model as reference like vanilla DPO, which occupies redundant GPU memory. Instead, we freeze the base model as reference and attach LoRA with an environment manager to flexibly switch between reference and action modes. Benefit from our dataset and techniques, PhyGDPO boosts the physical plausibility of the base T2V model and yields better results than the closed-source models OpenAI Sora2 [7] and Google Veo3.1 [20] on some challenging actions, as shown in Fig. 1, 4, 5, and 6.
In a nutshell, our contributions can be summarized as follows:
@@
  • We formulate a principled DPO framework, PhyGDPO, based on the groupwise Plackett-Luce probabilistic model to capture holistic physics advantage signal for physically consistent text-to-video generation.
@@
  • We design PGR to guide data sampling and preference optimization to focus on challenging physics. We propose LoRA-SR scheme to reduce GPU memory occupancy for more efficient and stable DPO training.
@@ 2.1 Physics-Augmented Video Data Construction
Due to the scarcity of text-video data pairs with rich physics interactions and phenomena, we design a Physics-Augmented video data construction Pipeline (PhyAugPipe). As shown in Fig. 2, we first use a VLM, Qwen2.5-72B-Instruct [69], equipped with our chain-of-thought (CoT) reasoning in (b) to derive physics-rich videos in (c) by filtering from a large T2V data pool in (a). Then in (d), we cluster the filtered data according to their action categories and sample the data with the rewards of a physics-aware VLM, VideoCon-Physics [4], in (e).
@@
Figure 3: Overview of PhyGDPO. (a) The LoRA-Switch Reference scheme can flexibly toggle between the reference and action modes for DPO advantage computation to save the GPU memory and increase the training stability. (b) our DPO framework is based on the groupwise Plackett-Luce (PL) probabilistic model and adopts a physics-aware VLM, VideoCon-Physics [4], to reward the DPO training. The winning case is the real-world video because it always captures correct physics as a guarantee.
Data Sampling with Physics Rewarding. As different action categories vary significantly in physical complexity thus leading to different generation difficulties for the T2V model, we propose to balance the training data accordingly. In particular, we first evaluate how well the T2V model performs on each action category. Based on the performance distribution, we adjust the sampling ratio by allocating more samples to the categories where the model performs poorly. As shown in Fig. 2 (d), we first select the top-$n_{c}$ samples with the highest semantics matching scores within each category as its representatives. Then we employ a physics-aware VLM, VideoCon-Physics [4], to evaluate each video representative. VideoCon-Physics outputs a semantics adherence score and a physics commonsense score $\in[0,1]$ to measure the physics plaucibility of the video. We average the two scores to obtain the overall score for each video. Then we compute the mean score of all representatives within each category to obtain the category score. In this way, we derive the score distribution histogram $\mathcal{S}_{f}$. Then we sample the data according to the difficulty of each action category. More specifically, we define the difficulty of the $k$-th action category as $d_{k}=1-\mathcal{S}_{f}(k)$ and assign a sampling weight following an exponential form as $r_{k}=\text{exp}(\tau d_{k})$, where $\tau$ is a hyperparameter that controls how strongly the sampling favors difficult action categories. Given the total sampling budget $N$, the number of data pairs sampled from the $k$-th action category is $\mathcal{H}_{r}(k)=\min\bigg(\mathcal{H}_{f}(k),\,N\cdot\frac{r_{k}}{\sum_{j=1}^{K_{a}}r_{j}}\bigg),$ where $\mathcal{H}_{r}$ denotes the distribution histogram after data sampling with physics rewarding. This sampling strategy allocates more data to challenging actions, encouraging the model to learn more complex physics during training.
Please note that our PhyAugPipe is a filtering-based data pipeline. The VLMs are used to select physics-rich scenes and hard actions from the real-world videos, which always follow the physical laws. Therefore, the VLM rewarding only impacts the physics richness rather than correctness. We use the original human-written prompts instead of the VLM rewritten ones to avoid the potential bias or errors of VLMs.
@@ 2.2 Physics-Aware Groupwise Direct Preference Optimization
Existing DPO algorithms are based on the Bradley–Terry (BT) probabilistic model, which compares pairwise data samples, showing limitations in capturing global and holistic preference signals and aligning with human feedback. Besides, vanilla DPO uses generated video as winning case but the physical realism of the generated video is limited and some challenging physics actions are difficult to generate. To address these issues, we formulate a Physics-Aware Groupwise Direct Preference Optimization (PhyGDPO), as illustrated in Fig. 3.
Groupwise Probability. We denote the reward as $r(c,x_{0})$ with the generation $x_{0}$ and condition $c$. Different from the normal DPO, we start from the groupwise Plackett-Luce (PL) probabilistic model, as shown in Fig. 3 (b). We adopt the real video as the winning case $x_{0}^{w}$ because it always follows physical laws and a set of generated videos as the losing cases $\mathcal{G}^{l}(c)=\{x_{0}^{l_{1}},\ldots,x_{0}^{l_{m}}\}$. The preference probability of the PL model is formulated as
@@
$\normalsize\mathcal{L}_{\mathrm{PL}}(\phi)=-\,\mathbb{E}_{c,\,\mathcal{G}^{l}(c)}\!\Bigg[\log\frac{\exp\!\big(r_{\phi}(c,x^{w}_{0})\big)}{\exp\!\big(r_{\phi}(c,x^{w}_{0})\big)+\sum_{j=1}^{m}\exp\!\big(r_{\phi}(c,x^{l_{j}}_{0})\big)}\Bigg].$ (2)
@@
$\normalsize f_{\theta}(x_{0},c)=\log\frac{p_{\theta}(x_{0}|c)}{p_{\psi}(x_{0}|c)}=\sum_{k=1}^{T}\log\frac{p_{\theta}(x_{t_{k-1}}\mid x_{t_{k}},c)}{p_{\psi}(x_{t_{k-1}}\mid x_{t_{k}},c)}\;\triangleq\;\sum_{k=1}^{T}\Delta_{k}.$ (5)
@@
Please refer to the supplementary for the detailed proof process of this inequality. Subsequently, we can further formulate the upper bound of $\mathcal{L}_{\text{GDPO}}(\theta)$ via Ineq. (7) into a single sample within each group for efficient training as
@@
$=\mathbb{E}_{c,\mathcal{G}^{l}(c),k,j}\Big[-m\gamma_{j}\log\sigma\big(\alpha_{j}\beta T(\Delta_{k}^{w}-\Delta_{k}^{l_{j}})\big)\Big].$
@@
Physics-Guided Rewarding. To further improve physics preference optimization, we design a Physics-Guided Rewarding (PGR) to direct DPO training to focus on challenging physics cases. More specifically, we parameterize $\gamma_{j}$ and $\alpha_{j}$ in Eq. (8) by the normalized semantics adherence score $s^{sa}_{j}\in[0,1]$ and physics commonsense score $s^{pc}_{j}\in[0,1]$ predicted by a physics-aware VLM, VideoCon-Physics, as
@@
$\alpha_{j}$ $=\alpha_{\text{min}}+(1-\alpha_{\text{min}})\cdot\sigma\big(\kappa_{\alpha}(v_{j}-b_{\alpha})\big),$
where $v_{j}$ measures the physics difficulty and mainly modulates $\gamma_{j}$ and $\alpha_{j}$, $\alpha_{\text{min}}$ sets a lower bound to avoid vanishing gradients, and the sigmoid function $\sigma(\cdot)$ ensures a bounded and smooth adjustment of $\gamma_{j}$, stabilizing the optimization. $\alpha_{j}$ adaptively adjusts the sharpness of the preference comparison, while $\gamma_{j}$ amplifies the physics-guided reward for samples with lower physical plausibility. They are controlled by the hyperparameters $\lambda$, $\kappa_{\alpha,\gamma}$, $b_{\alpha,\gamma}$, and $\alpha_{\text{min}}$. Our designed physics rewards $(\alpha_{j},\gamma_{j})$ balances the training stability with adaptivity, allowing physics-violating samples to exert stronger influence during optimization. Please note that the VLM rewarding only affects the importance of samples while the correctness of physical supervision is guaranteed by the core objective, i.e., the DPO advantage of winning case (real-world video capturing correct physics) over losing case.
@@
$\log\frac{p_{\theta}(x_{t-h}\mid x_{t},c)}{p_{\psi}(x_{t-h}\mid x_{t},c)}$ $\approx-\frac{h^{2}}{2\varepsilon}\Big(\ell_{\theta}(x_{t},t)-\ell_{\psi}(x_{t},t)\Big),$ (12)
@@
$\normalsize\mathcal{L}=\mathbb{E}_{c,\mathcal{G}^{l}(c),k,j}\Big[-m\gamma_{j}\log\sigma\!\big(-\alpha_{j}\beta T[(\ell_{\theta}^{w}-\ell_{\psi}^{w})-(\ell_{\theta}^{l_{j}}-\ell_{\psi}^{l_{j}})]\big)\Big],\vskip 2.84526pt$ (13)
where any global constant scale factor such as $h^{2}/2\varepsilon$ in Eq. (12) can be absorbed into the hyperparameter $\beta$. For simplicity, we denote $\ell_{\theta}^{w}=\ell_{\theta}(x_{t_{k}}^{w}\mid t_{k})$, $\ell_{\psi}^{w}=\ell_{\psi}(x_{t_{k}}^{w}\mid t_{k})$, $\ell_{\theta}^{l_{j}}=\ell_{\theta}(x_{t_{k}}^{l_{j}}\mid t_{k})$, and $\ell_{\psi}^{l_{j}}=\ell_{\psi}(x_{t_{k}}^{l_{j}}\mid t_{k})$ in Eq. (13).
LoRA-Switch Reference. Previous DPO methods usually copy the full model and fix it as the reference, which occupies redundant GPU memory, degrades computational efficiency, and hinders the scalability of model size. Plus, this full-copy strategy often leads to unstable and less effective DPO training because it updates the entire set of the weights and may cause the action model to quickly and dramatically deviate from the reference model.
@@
To address this issue, we design a LoRA-Switch Reference (LoRA-SR) scheme. As shown in Fig. 3 (b), we freeze the backbone as the reference model $\psi$ and attach trainable LoRA modules to $\psi$ as the action model $\theta$. Specifically, we attach LoRA to the query, key, value, and output linear projection layers of each self-attention in the Transformer backbone of the T2V model and implement an environment manager as the switch to flexibly toggle between the reference and action modes. Given the input tokens of the linear projection layer where the LoRA is attached to as $\mathbf{X}\in\mathbb{R}^{B\times L\times C_{in}}$, the output tokens $\mathbf{Y}\in\mathbb{R}^{B\times L\times C_{in}}$ of our LoRA-SR scheme can be formulated as
$\small\mathbf{Y}=\mathbf{X}(\mathbf{W}+\mathbf{1}_{action}\cdot\Delta\mathbf{W})^{\top}=\mathbf{X}(\mathbf{W}+\mathbf{1}_{action}\cdot\frac{\alpha}{r}\mathbf{B}\mathbf{A})^{\top},$ (14)
where $\mathbf{W}\in\mathbb{R}^{C_{out}\times C_{in}}$ is the frozen weight of the linear layer, $\mathbf{1}_{action}\in\{0,1\}$ is a binary indicator governed by the environment manager to switch between the reference ($\mathbf{1}_{action}=0$) and action ($\mathbf{1}_{action}=1$) modes, $\alpha$ is the scaling factor, $r$ is the rank, and $\mathbf{B}\in\mathbb{R}^{C_{out}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times C_{in}}$ are the learnable LoRA parameters. Our LoRA-SR enables the reference and trainable models to share the same heavy backbone while flexibly switching the lightweight LoRA parameters for the action and reference models. Thus, we can compute the DPO advantage without storing or loading an extra full model. This significantly increases computational efficiency, improves the scalability of model size, and prevents the model from drifting too far from reference, thus stabilizing the training process.